# bibliography.bib Post-Training Network Compression for 3D Medical Image Segmentation: Reducing Computational Efforts via Tucker Decomposition

### v1

@@ 2.2 Dataset
The TS package provides access to two models: The default model for 1.5mm resolution and a fast version for 3mm resolution. It is important to note that, while being marketed as a single-model interface, the actual model consists of a conglomerate of five independent U-nets, all trained separately on a different set of label classes, i.e., one separate model for the skeleton, cardiovascular system, gastrointestinal tract, muscles, and other organs.
@@ 4 Discussion
Another established method in weight factorization, CP-decomposition [lebedev2015speeding], factorizes each dimension of the target kernel. In our scenario, this approach would necessitate the addition of five layers for each decomposed convolution. We opted against this measure because, as previously discussed, extending the neural network sequence raises the minimum runtime lower bound. Adding five layers rather than three reduces the performance gains on modern GPU hardware. Furthermore, studies have reported instability issues with CP-decomposition [de2008tensor, lebedev2015speeding, yong2016tucker]. To the best of our knowledge, our study is the first to explore network compression with Tucker decomposition in medical image segmentation and its implications for clinical practice.
@@
As presented in the results section, the inference speedups did not directly correspond to the theoretical reduction in FLOPs. The GTX series cards are not equipped with tensor cores, therefore exhibiting larger performance gains using 16-bit precision. The larger variances and extended inference times can be attributed to the overhead associated with dynamic conversion from 32-bit to 16-bit precision, coupled with the absence of dedicated hardware to execute accelerated algorithmic operations efficiently. For practitioners not limited by memory constraints, we recommend avoiding using 16-bit precision casting on the GTX series GPUs.
Another observation was that for large DF, the execution time of the compressed model can be higher than the full original model. While in theory, the number of FLOPs was reduced, the sequential nature of replacing one layer with a series of three layers came with additional overhead. For most consumer cards, there was still a notable speedup with runtimes twice as fast as the original model. High-performance data center GPUs, such as the NVIDIA A100, possess a sufficient number of CUDA cores to process the original models with a speed comparable to that of their compressed counterparts.

### v2

@@ 2.2 Dataset
The TS package provides access to two models: The default model for 1.5mm resolution and a fast version for 3mm resolution. Although marketed as a single-model interface, the 1.5mm model consists of five separate U-Nets, trained on different label sets, i.e., one separate model for the skeleton, cardiovascular system, gastrointestinal tract, muscles, and other organs.
@@ 4 Discussion
Another established method in weight factorization, CP-decomposition [lebedev2015speeding], factorizes each dimension of the target kernel. In our scenario, this approach would necessitate the addition of five layers for each decomposed convolution. We opted against this measure because, as previously discussed, extending the neural network sequence raises the minimum runtime lower bound. Adding five layers rather than three reduces the performance gains on modern GPU hardware. Furthermore, studies have reported instability issues with CP-decomposition [de2008tensor, lebedev2015speeding, yong2016tucker]. Other studies have explored the integration of weight quantization into CP-decomposition [cherniuk2023quantization] and Tucker decomposition [liu2024towards]. Hierarchical Tucker decomposition was proposed to address the challenges in compressing recurrent neural networks [yin2020compressing]. Moreover, [cheng2020novel] tackled the problem of optimal tensor rank selection using reinforcement learning, whereas [idelbayev2020low] formulated this as a discrete-continuous optimization task. [bhalgaonkar2024model] highlighted the priority of accuracy over speed when using network compression in medical diagnostics. Closely related to our domain, [ashtari2021low] utilized a Tucker-inspired low-rank convolution kernel as a regularization strategy for the training of brain tumor segmentation models. To the best of our knowledge, our study is the first to explore network compression with Tucker decomposition in medical image segmentation and its implications for clinical practice.
@@
The results indicate that inference speedups did not match the theoretical FLOP reductions. The GTX series cards, lacking tensor cores, do not benefit from performance gains using 16-bit precision due to the absence of specialized hardware for efficient accelerated computation. For practitioners without memory constraints, we recommend avoiding using 16-bit precision on GTX series GPUs.
Another observation was that the execution time of the compressed model can be higher than the full original model. While in theory, the number of FLOPs was reduced, the sequential nature of replacing one layer with a series of three layers came with additional overhead. For most consumer cards, there was still a notable speedup with runtimes twice as fast as the original model. High-performance data center GPUs, such as the NVIDIA A100, possess a sufficient number of CUDA cores to process the original models with a speed comparable to that of their compressed counterparts.